OpenStudy (anonymous):

Using substitution cosx=u integrate sin3x.cos4x and show it is equal to 17cos7x−15cos5x+c

OpenStudy (anonymous):

Sorry the 17cos7x−15cos5x+c is wrong it's meant to be 1/7cos^7 x -1/5cos^5 x

OpenStudy (anonymous):

lol. that changes everything :) let u = cos x ---> du = -sinx dx

OpenStudy (anonymous):

remember that sin^2(x) = 1 - cos^2(x). you must use this identity when doing trig integration. so we have (1 - cos^2(x))*cos^4(x)*sinx dx = (1 - u^2)*u^4 du

OpenStudy (anonymous):

Also the top bit is sin^3 x cos^4 x, not sure why it didn't go to a power when writing the equation

OpenStudy (anonymous):

what i said times - 1 since du = - sinx

OpenStudy (anonymous):

so (u^2 - 1)u^4 du

OpenStudy (anonymous):

should i keep going?

OpenStudy (anonymous):

Yes I understand so far thanks

OpenStudy (anonymous):

not sure if if that means i should or not :P

OpenStudy (anonymous):

yes keep going please :p

OpenStudy (anonymous):

= integral of u^6 - u^4 = (1/7)u^7 - (1/5)u^5 + C plug back u = cos x = (1/7)cos^7(x) - (1/5)cos^5(x) + C

OpenStudy (raden):

better we use the identity : sinAcosB = 1/2 (sin(A+B)+sin(A-B)) so, sin3x cos4x = 1/2 (sin(3x+4x) + sin(3x-4x) = 1/2 (sin7x - sinx) now, see ur integral becomes